RAS2TIFF(1) RAS2TIFF(1) NAME ras2tiff−createaTIFFfilefromaSunrasterfile SYNOPSIS ras2tiff[options]input.rasoutput.tif DESCRIPTION ras2tiffconvertsafileintheSunrasterfileformattoTIFF.Bydefault,theTIFFimageiscreated withdatasamplespacked(PlanarConfiguration=1),compressedwiththeLempel-Ziv&Welch algorithm(Compression=5),andwitheachstripnomorethan8kilobytes.Thesecharacteristics canoverriden,orexplicitlyspecifiedwiththeoptionsdescribedbelow. AnycolormapinformationintherasterfileiscarriedovertotheTIFFfilebyincludingaColormap tagintheoutputfile.Iftherasterfilehasacolormap,thePhotometricInterpretationtagissetto 3(palette);otherwiseitissetto2(RGB)ifthedepthis24or1(min-is-black)ifthedepthisnot 24. OPTIONS −cSpecifyacompressionschemetousewhenwritingimagedata:−cnonefornocompression,-cpackbitsforthePackBitscompressionalgorithm,-cjpegforthebaselineJPEG compressionalgorithm,-czipfortheDeflatecompressionalgorithm,and−clzwfor Lempel-Ziv&Welch(thedefault). −rWritedatawithaspecifiednumberofrowsperstrip;bydefaultthenumberofrows/strip isselectedsothateachstripisapproximately8kilobytes. BUGS Doesnothandleallpossiblerasterfiles.Inparticular,ras2tiffdoesnothandlerun-lengthencoded images. SEEALSO pal2rgb(1),tiffinfo(1),tiffcp(1),tiffmedian(1),libtiff(3) October15,1995 1